# Dynamic Learning Rate for Deep Reinforcement Learning: A Bandit Approach

### v1

@@ B.1 Evaluation Terminology
  • Jumpstart performance: The performance at the initial stages of training, starting from a policy with randomized parameters $\theta$. In Figure 7, Algorithm B exhibits better jumpstart performance but ultimately achieves lower final performance than Algorithm A. A lower jumpstart performance can result from factors such as a lower learning rate, although this work demonstrates that this does not necessarily lead to worse final performance.
Figure 7: Performance curves of two RL algorithms (adapted from taylor2009transfer [36]).
B.2 Further Experimental Details
@@
    $\hskip 25.03856pt\mathcal{K}_{\text{middle}}(3)=\bigl{\{}3.125\times 10^{-5},6.25\times 10^{-5},1.25\times 10^{-4}\bigr{\}}$
    $\hskip 84.50471pt\mathcal{K}_{\text{highest}}(3)=\bigl{\{}6.25\times 10^{-5},1.25\times 10^{-4},2.5\times 10^{-4}\bigr{\}}$
@@
Section 5.3 – RMSProp Optimizer and More Environments.
  • Optimizer: RMSProp for baselines
@@
  • LRRL set of learning rates: $\mathcal{K}=\bigl{\{}1.5625\times 10^{-5},3.125\times 10^{-5},6.25\times 10^{-5}\bigr{\}}$
Section A.2 – A Comparison between Multi-Armed Bandit Algorithms.
@@
B.3 Hyperparameters
In this section, we outline the hyperparameters used in the experiments. The optimizers from the Optax library [11] are employed alongside the JAX [8] implementation of the DQN algorithm [24, 25], as provided by the Dopamine framework [9].
Hyperparameter Setting
@@ B.3 Hyperparameters
Exploration decay period 250000 steps
Environment steps per iteration 250000 steps
@@
Eps 1.5e-4
RMSProp hyperparameters
@@
Eps 1.5e-4
Table 4: Hyperparameters

### v2

@@ Appendix C Evaluation Metrics
  • Jumpstart performance: The performance at the initial stages of training, starting from a policy with randomized parameters $\theta$. In Figure 12, Algorithm A exhibits better jumpstart performance but ultimately achieves lower final performance than Algorithm B. A lower jumpstart performance can result from factors such as a lower learning rate, although this work demonstrates that this does not necessarily lead to worse final performance.
Figure 12: Performance curves of two RL algorithms (adapted from 40).
Appendix D Implementation Details
@@
    $\mathcal{K}_{\text{middle}}(3)$ =$=\hskip 65.44142pt\bigl{\{}3.125\times 10^{-5},6.25\times 10^{-5},1.25\times 10^{-4}\bigr{\}}$
    $\mathcal{K}_{\text{highest}}(3)$ =$=\hskip 125.19194pt\bigl{\{}6.25\times 10^{-5},1.25\times 10^{-4},2.5\times 10^{-4}\bigr{\}}$
@@
Section 4.3– Comparing RMSProp and Adam Optimizers
  • Optimizer: RMSProp for DQN baselines
@@
  • LRR-DQN set of learning rates: $\mathcal{K}=\bigl{\{}1.5625\times 10^{-5},3.125\times 10^{-5},6.25\times 10^{-5}\bigr{\}}$
Section 4.4– LRRL with a Distributional RL Algorithm.
  • Optimizer: Adam
  • Bandit step-size: $\alpha=0.2$
  • LRRL-IQN set of learning rates: $\mathcal{K}=\bigl{\{}2.5\times 10^{-5},5\times 10^{-5},1\times 10^{-4}\bigr{\}}$
Section B.2 – A Comparison between Multi-Armed Bandit Algorithms.
@@
Hyperparameters of the optimizers and deep RL algorithms.
For reproductibility, the list of hyperparameters of the optimizers and deep RL algorithms used for our experiments is provided in Table 1. The optimizers from the Optax library [12] are employed alongside the JAX [8] implementation of the DQN [25, 26] and IQN [11] algorithms, as provided by the Dopamine framework [9].
DQN and IQN hyperparameters Value
@@
Exploration decay period (steps) 250 000
Environment steps per iteration (steps) 250 000
@@
Eps (DQN) 1.5$\times$10-4
Eps (IQN) 3.125$\times$10-4
RMSProp hyperparameters (DQN)
@@
Eps 1.5$\times$10-4
Table 1: Hyperparameters used in the experiments.